# Entanglement harvesting and curvature of entanglement: A modular operator approach

### v1

@@ VII.1 Tomita–Takesaki Modular Operators
Consider a $\bm{C^{*}}$-algebra $\mathcal{B(H)}=\{A\}$ of bounded linear operators on a Hilbert space, $A:\mathcal{H}\rightarrow\mathcal{H}$. Let $\mathcal{C}$ be a subset of $\mathcal{B(H)}$. An operator $A\in\mathcal{B(H)}$ belongs to the commutant $\mathcal{C}^{\prime}$ of the set $\mathcal{C}$ $\iff$ $AC=CA,\,\,\,\forall C\in\mathcal{C}$. A von Neumann algebra $\mathcal{A}$ is a unital $C^{*}$-subalgebra of $\mathcal{B(H)}$ such that $\mathcal{A}^{\prime\prime}=\mathcal{A}$. Von Neumann proved that this bicommutant definition is equivalent to the algebra $\mathcal{A}$ being closed with respect to the weak topology on $\mathcal{B(H)}$. A von Neumann algebra in standard form is one where there exists an element $|\Omega\rangle\in\mathcal{H}$ which is both cyclic (operating on $|\Omega\rangle$ with elements in $\mathcal{A}$ can generate a space dense in $\mathcal{H}$) and separating (if $A|\Omega\rangle=0$, then $A=0$).
@@ VII.2.2 Qubits
We consider the special case of modular conjugation operator $J_{AB}$ acting on product qubit states in $\mathcal{H}_{A}\otimes\mathcal{H}_{B}\cong\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, defined in the computational basis as
@@
$\sigma_{Y}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},\quad Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$. (88)
Let us compute the action of $J_{\mathrm{QIT}}$ on a generic product state
@@
Clearly, the two results differ by a SWAP and additional local sign flips. This is captured precisely by local Pauli-$Z$ operators. We therefore have:
$J_{AB}=(Z\otimes Z)\,\mathrm{SWAP}\circ J_{\mathrm{QIT}},$ (91)
@@
All other terms such as $\langle\Psi^{(0)}_{f}|J_{AB}\otimes\mathbb{I}_{\phi}|\Psi^{(0)}_{f}\rangle$ vanish due to orthogonality, and $\langle\Psi^{(2)}_{f}|J_{AB}|\Psi^{(2)}_{f}\rangle\sim\lambda^{4}$. First,

### v2

@@ VII.1 Tomita–Takesaki Modular Operators
Consider a $\bm{C^{*}}$-algebra $\mathcal{B(H)}=\{A\}$ of bounded linear operators on a Hilbert space, $A:\mathcal{H}\rightarrow\mathcal{H}$. Let $\mathcal{C}$ be a subset of $\mathcal{B(H)}$. An operator $A\in\mathcal{B(H)}$ belongs to the commutant $\mathcal{C}^{\prime}$ of the set $\mathcal{C}$ $\iff$ $AC=CA,\,\,\,\forall C\in\mathcal{C}$. A von Neumann algebra $\mathcal{A}$ is a unital $\bm{C^{*}}$-subalgebra of $\mathcal{B(H)}$ such that $\mathcal{A}^{\prime\prime}=\mathcal{A}$. Von Neumann proved that this bicommutant definition is equivalent to the algebra $\mathcal{A}$ being closed with respect to the weak topology on $\mathcal{B(H)}$. A von Neumann algebra in standard form is one where there exists an element $|\Omega\rangle\in\mathcal{H}$ which is both cyclic (operating on $|\Omega\rangle$ with elements in $\mathcal{A}$ can generate a space dense in $\mathcal{H}$) and separating (if $A|\Omega\rangle=0$, then $A=0$).
@@ VII.2.2 Qubits
We consider the special case of the modular conjugation operator $J_{AB}$ acting on product qubit states in $\mathcal{H}_{A}\otimes\mathcal{H}_{B}\cong\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, defined in the computational basis as
@@
$\sigma_{y}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$ (94)
and $K$ denotes complex conjugation in the computational basis. Let us compute the action of $J_{\mathrm{QIT}}$ on a generic product state
@@
Clearly, the two results differ by a SWAP and additional local sign flips. This is captured precisely by local Pauli-$\sigma_{z}$ operators
$\sigma_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$ (97)
such that (up to global phase)
$J_{AB}=(\sigma_{z}\otimes\sigma_{z})\,\mathrm{SWAP}\circ J_{\mathrm{QIT}},$ (98)
where
$\mathrm{SWAP}\begin{pmatrix}\alpha\\ \beta\end{pmatrix}\otimes\begin{pmatrix}\gamma\\ \delta\end{pmatrix}=\begin{pmatrix}\gamma\\ \delta\end{pmatrix}\otimes\begin{pmatrix}\alpha\\ \beta\end{pmatrix}$ (99)
@@
All other terms such as $\langle\Psi^{(0)}_{f}|J_{AB}\otimes\mathbb{I}_{\phi}|\Psi^{(0)}_{f}\rangle$ vanish due to orthogonality, and $\langle\Psi^{(2)}_{f}|J_{AB}\otimes\mathbb{I}_{\phi}|\Psi^{(2)}_{f}\rangle\sim\lambda^{4}$. First,
@@
Acknowledgements.
The author thanks an anonymous referee for several insightful comments that improved the clarity and scope of this work.